OpenStudy (anonymous):

Solve the following formula for R: C=2πR Choose ALL equivalent variations of your answer. I know one equivalent variation is r=c/2π but I'm not too sure which other ones are equivalent variations.

OpenStudy (anonymous):

Here are the options

OpenStudy (unklerhaukus):

say we have an equation like this xy=z we can divide both sides by y xy/y=z/y x=z/y

OpenStudy (anonymous):

@UnkleRhaukus Yes, I divided both sides by 2π and got r=c/2π. But the question says to find all variations of that answer, that is the part I'm stuck on, finding which other one is a variation of my answer.

OpenStudy (unklerhaukus):

x=z/y is the same as x=1/y * z

OpenStudy (anonymous):

@UnkleRhaukus Oh okay, thank you!